Shrieves Casting Company is considering adding a new line to its product mix, and the capital budgeting analysis is being conducted by Sidney Johnson, a recently graduated MBA. The cost of new machinery for the new product line would be $644,000. The machinery has economic life of eight years, and has no salvage value. Assume that depreciation is straight-line to zero over the life of the machine. The new line would generate incremental sales of 70,000 units per year at variable cost per unit of $21 and fixed cost of $725,000 per year. Each unit can be sold for $37 in the first year. The sale price and cost are both expected to remain the same. The firm’s tax rate is 35 percent, and the rate of return required for this type of investment is 15 percent.

  1. Calculate the base-case cash flow and NPV.
  2. Calculate the accounting break-even point.
  3. What is the sensitivity of NPV to changes in the sales quantity? What is the effect on NPV of a 500-unit decrease in projected sales.
  4.   Suppose the projections given for price, quantity, variable costs, and fixed costs are all

        accurate to within +/-10 percent. Calculate the best-case and worst-case NPV.
Answer #1

a. Depreciation expense per year=644000/7=92000

Cash flow would be= Revenue-Variable cost-Fixed Cost-Tax

Below is the calculation of cash flow and NPV

Year Sales Unit Intial Cost Revenue Variable Cost Fixed Cost Depreciatio Expense (Intial Cost/7) Tax (35%*(Revenue-Variable Cost-Fixed Cost-Depreciation) Cash Flow Discounted Cash Flow (Cash Flow/1.15^year)
0 0 (644,000.00)                      -                        -                     -                                                               -                                                            -          (644,000.00)                          (644,000.00)
1 70000                     -   2,590,000.00 1,470,000.00 725,000.00                                               92,000.00                                          106,050.00          288,950.00                            251,260.87
2 70000                     -   2,590,000.00 1,470,000.00 725,000.00                                               92,000.00                                          106,050.00          288,950.00                            218,487.71
3 70000                     -   2,590,000.00 1,470,000.00 725,000.00                                               92,000.00                                          106,050.00          288,950.00                            189,989.32
4 70000                     -   2,590,000.00 1,470,000.00 725,000.00                                               92,000.00                                          106,050.00          288,950.00                            165,208.10
5 70000                     -   2,590,000.00 1,470,000.00 725,000.00                                               92,000.00                                          106,050.00          288,950.00                            143,659.22
6 70000                     -   2,590,000.00 1,470,000.00 725,000.00                                               92,000.00                                          106,050.00          288,950.00                            124,921.06
7 70000                     -   2,590,000.00 1,470,000.00 725,000.00                                               92,000.00                                          106,050.00          288,950.00                            108,627.01
NPV (Total of discounted cash flow)                            558,153.28

b. We can see that initial expense are being recovered in the  3rd year.

Accounting break even period= 2+(644000-2*288950)/288950=2.22 year

c.

If the units are decreases by 500 then NPV decreased to $536519.1

Year Sales Unit Intial Cost Revenue Variable Cost Fixed Cost Depreciatio Expense (Intial Cost/7) Tax (35%*(Revenue-Variable Cost-Fixed Cost-Depreciation) Cash Flow Discounted Cash Flow (Cash Flow/1.15^year)
0 0 (644,000.00)                      -                        -                     -                                                               -                                                            -          (644,000.00)                          (644,000.00)
1 69500                     -   2,571,500.00 1,459,500.00 725,000.00                                               92,000.00                                          103,250.00          283,750.00                            246,739.13
2 69500                     -   2,571,500.00 1,459,500.00 725,000.00                                               92,000.00                                          103,250.00          283,750.00                            214,555.77
3 69500                     -   2,571,500.00 1,459,500.00 725,000.00                                               92,000.00                                          103,250.00          283,750.00                            186,570.23
4 69500                     -   2,571,500.00 1,459,500.00 725,000.00                                               92,000.00                                          103,250.00          283,750.00                            162,234.98
5 69500                     -   2,571,500.00 1,459,500.00 725,000.00                                               92,000.00                                          103,250.00          283,750.00                            141,073.90
6 69500                     -   2,571,500.00 1,459,500.00 725,000.00                                               92,000.00                                          103,250.00          283,750.00                            122,672.96
7 69500                     -   2,571,500.00 1,459,500.00 725,000.00                                               92,000.00                                          103,250.00          283,750.00                            106,672.14
NPV (Total of discounted cash flow)                            536,519.10

So, sensitivity= ( 558158.23-536519.1)/558158.23=3.88%

d.

For Best case scenario price will be 37*1.1=$40.7/unit, quantity=70000*1.1=77000 and variable cost is 21*(1-10%)=18.9 and fixed cost=725000*(1-10%)=$6525000, So, NPV calculation is given below:

Year Sales Unit Intial Cost Revenue Variable Cost Fixed Cost Depreciatio Expense (Intial Cost/7) Tax (35%*(Revenue-Variable Cost-Fixed Cost-Depreciation) Cash Flow Discounted Cash Flow (Cash Flow/1.15^year)
0 0 (644,000.00)                      -                        -                     -                                                               -                                                            -          (644,000.00)                          (644,000.00)
1 77000                     -   3,133,900.00 1,455,300.00 652,500.00                                               92,000.00                                          326,935.00          699,165.00                            607,969.57
2 77000                     -   3,133,900.00 1,455,300.00 652,500.00                                               92,000.00                                          326,935.00          699,165.00                            528,669.19
3 77000                     -   3,133,900.00 1,455,300.00 652,500.00                                               92,000.00                                          326,935.00          699,165.00                            459,712.34
4 77000                     -   3,133,900.00 1,455,300.00 652,500.00                                               92,000.00                                          326,935.00          699,165.00                            399,749.86
5 77000                     -   3,133,900.00 1,455,300.00 652,500.00                                               92,000.00                                          326,935.00          699,165.00                            347,608.57
6 77000                     -   3,133,900.00 1,455,300.00 652,500.00                                               92,000.00                                          326,935.00          699,165.00                            302,268.32
7 77000                     -   3,133,900.00 1,455,300.00 652,500.00                                               92,000.00                                          326,935.00          699,165.00                            262,842.02
NPV (Total of discounted cash flow)                         2,264,819.86

Hence for best case scenario NPV=$2264819.86

For worst case scenario, units=70000*0.9=63000, price=37*0.9=33.3, variable cost=21*1.1=23.1 and fixed cost=725000*1.1=797500

NPV calculation is given below:

Hence for worst case scenario NPV is -$928926.34